Find the evolute of the four cusped hypocycloid
x=a cos^3θ y=a sin^3θ
1
Expert's answer
2020-11-26T13:40:42-0500

Find the derivatives of "x" and "y" with respect to the parameter "\theta :"


"x'=-3a\sin \theta\cos^2\theta,"


"y'=3a\sin^2 \theta\cos\theta,"

"x''=-3a\cos^3\theta+6a\sin^2 \theta\cos\theta"


"y''=-3a\sin^3 \theta+6a\sin \theta\cos^2\theta"

To calculate the coordinates of the center of curvature "(\xi,\eta),"  we use the formulas:


"\xi=x-y'\dfrac{(x')^2+(y')^2}{x'y''-x''y'}"

"\eta=y+x'\dfrac{(x')^2+(y')^2}{x'y''-x''y'}"

"(x')^2+(y')^2=9a^2 \sin^2\theta\cos^4\theta+9a^2 \sin^4\theta\cos^2\theta="

"=9a^2 \sin^2\theta\cos^2\theta"

"x'y''-x''y'=9a^2\sin^4 \theta\cos^2\theta-18a^2\sin^2 \theta\cos^4\theta"

"+9a^2\sin^2 \theta\cos^4\theta-18a^2\sin^4 \theta\cos^2\theta="

"=9a^2\sin^2 \theta\cos^2\theta-18a^2\sin^2 \theta\cos^2\theta"

"=-9a^2\sin^2 \theta\cos^2\theta"


"\dfrac{(x')^2+(y')^2}{x'y''-x''y'}=-1"

"\xi=a\cos^3\theta+3a\sin^2\theta \cos\theta"

"\eta=a\sin^3\theta+3a\sin\theta \cos^2\theta"

Consequently, the evolute of the four cusped hypocycloid is described by the following parametric equations:


"\xi=a\cos^3\theta+3a\sin^2\theta \cos\theta"

"\eta=a\sin^3\theta+3a\sin\theta \cos^2\theta"


Thus the evolute of an astroid is an astroid .